OpenStudy (anonymous):

I'm doing a worksheet for algebra II, and I need some help with a concept. Here's the question and my work: Create a new function, h(x). Then assign any number to x. Using complete sentences, explain whether f(h(x)) and h(f(x)) will always result in the same number. You will use the function f(x) that you created in problem number 2. f(x) = x^2 + 2 and h(x) 2x – 4. Let x equal 6. h(x) = 2(6) – 4 f(h(x)) = (2(6) – 4)^2 + 2 f(h(x)) = (12 – 4)^2 + 2 f(h(x)) = 8^2 + 2 f(h(x)) = 64 + 2 = 66 h(f(x)) = 2(6^2 + 2) – 4 h(f(x)) = 2(36 + 2) – 4 h(f(x)) = 76 – 4 h(f(x)) = 72

OpenStudy (anonymous):

So my question is - SHOULD those be getting the same answer?

OpenStudy (anonymous):

Is there something I need to change in my approach to get the same answers?

OpenStudy (anonymous):

@random231 Can you help again?

OpenStudy (anonymous):

@brandonE

OpenStudy (rock_mit182):

in general, \[f ° g \neq g°f\] . Remember that the notation f g means that the function g is applied first and then f is applied second.

OpenStudy (anonymous):

@rock_mit182 It's not supposed to be equal?

OpenStudy (rock_mit182):

in general is supposed, cause is no commutative ( like in sum,3+5=5+3) in this case the order matters

OpenStudy (rock_mit182):

besides their domains can differ

OpenStudy (anonymous):

Okay. Thank you :)

OpenStudy (anonymous):

Did I get the order right in applying f(x) first?

OpenStudy (anonymous):

@rock_mit182

OpenStudy (rock_mit182):

sure your work is clear

OpenStudy (rock_mit182):

f ° g = f (g(x)) ; g ° f = g (f(x) ;

OpenStudy (rock_mit182):

yeah good job ;)

OpenStudy (anonymous):

@random231 Can you help me once more w/ inverse functions? f(x) = x^2 + 2 when f(x) = 3.

OpenStudy (anonymous):

Thank you! Thanks for your help :)

random231 (random231):

ya sure!!

OpenStudy (anonymous):

@random231 x = y^2 + 2...do I substitute 3 for y?

OpenStudy (rock_mit182):

yeah but you halve to solve for y then

OpenStudy (rock_mit182):

have*

random231 (random231):

exactly! actually when u find the inverse of a function and write it as a function of x,JUST SUBSTITUTE Y WITH X.

OpenStudy (rock_mit182):

then plug the value 3

random231 (random231):

yup!

OpenStudy (anonymous):

Okay, could you write out what that would look like? I'm a little confused, sorry :( Well, not confused, but I need a visual.

random231 (random231):